OpenStudy (anonymous):

A circle has its center at (-2, 5) and a radius of 4 units. What is the equation of the circle?

OpenStudy (moongazer):

\[(x-h)^2 + (y-k)^2 = r^2 \] if c(h,k)

OpenStudy (moongazer):

then transfer to the Gen. Form: \[Ax^2 + By^2 + Cx +Dy +E = 0\]

OpenStudy (moongazer):

just substitute the values

OpenStudy (moongazer):

then solve :)

OpenStudy (anonymous):

Well these are the choices, can you help solve it

OpenStudy (anonymous):

first option